### v1

# Minimal plus-one generated curves and Briançon-type polynomials

Abstract.
We define the minimal plus-one generated curves and prove a result explaining why they are the closest relatives of the free curves, after the nearly free curves. Then we look at the projective closures of the general and of the special fibers of some Briançon-type polynomials constructed by E. Artal Bartolo, Pi. Cassou-Noguès and I. Luengo Velasco. They yield new examples of free, nearly free or minimal plus-one generated curves. This leads us to ask the natural question if such a result holds for any Briançon-type polynomials.
Key words and phrases:
plane curve, Milnor algebra, minimal resolution, Milnor and Tjurina number, Briançon-type polynomial
2010 Mathematics Subject Classification:
Primary 14H50; Secondary 13D02
1. Introduction
In this note we first define the minimal plus-one generated curves in Definition 1.1 and give a simple numerical characterization of them in Theorem 1.2. This characterization is similar to the known characterizations of free and nearly free curves, which are recalled in Corollaries 2.2 and 2.3. Then we construct new examples of free, nearly free and minimal plus-one generated curves starting with some Briançon-type polynomials constructed in [3].
Let us recall the definitions of all these objects. Let $S=\mathbb{C}[x,y,z]$ be the polynomial ring in three variables $x,y,z$ with complex coefficients, and let $C:f=0$ be a reduced curve of degree $d\geq 3$ in the complex projective plane $\mathbb{P}^{2}$. We denote by $J_{f}$ the Jacobian ideal of $f$, i.e., the homogeneous ideal in $S$ spanned by the partial derivatives $f_{x},f_{y},f_{z}$ of $f$, and by $M(f)=S/J_{f}$ the corresponding graded quotient ring, called the Jacobian (or Milnor) algebra of $f$. Consider the graded $S$-module of Jacobian syzygies of $f$ or, equivalently, the module of derivations killing $f$, namely
(1.1) $D_{0}(f)=\{\theta\in Der(S)\ :\ \theta(f)=0\}=\{\rho=(a,b,c)\in S^{3}\ :\ af_{x}+bf_{y}+cf_{z}=0\}.$
We say that $C:f=0$ is an $m$-syzygy curve if the module $D_{0}(f)$ is minimally generated by $m$ homogeneous syzygies, say $\rho_{1},\rho_{2},\ldots,\rho_{m}$, of degrees $d_{j}=\deg\rho_{j}$ ordered such that
$d_{1}\leq d_{2}\leq\ldots\leq d_{m}.$
In this note we assume that $C$ is not the union of $d$ lines passing through one point, hence $d_{1}\geq 1$. We call these degrees $(d_{1},\ldots,d_{m})$ the exponents of the curve $C$. The smallest degree $d_{1}$ is sometimes denoted by ${\rm mdr}(f)$ and is called the minimal degree of a Jacobian relation for $f$.
The curve $C$ is free when $m=2$, since then $D_{0}(f)$ is a free module of rank 2, see for instance [22, 23, 24, 25]. A free curve $C$ is also characterized by the condition $d_{1}+d_{2}=d-1$, and hence in this case $r=d_{1}<d/2$, see [24]. A nearly free curve is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$ and $d_{2}=d_{3}$. To simplify the notation, we say that $(d_{1},d_{2})$ are the exponents for the nearly free curve $C$. Finally, a plus-one generated curve is a 3-syzygy curve $C$ with exponents $(d_{1},d_{2},d_{3})$ such that $d_{1}+d_{2}=d$, see [1, 15]. We define the $\delta$-level of a plus-one generated curve by the formula
(1.2) $\delta L(C)=d_{3}-d_{2}\geq 0.$
Note that a plus-one generated curve $C$ is nearly free if and only if $\delta L(C)=0$.
Definition 1.1.
A plus-one generated curves $C$ satisfying $\delta L(C)=1$ is called a minimal plus-one generated curve, or an $MPOG$-curve for short.
These curves can be regarded as the closest neighbors of the free curves beyond the nearly free curves, as shown by our main result in this note.
Theorem 1.2.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$. Then $C$ is an $MPOG$-curve if and only if
$\tau(C)=\tau(d,r)_{max}-2.$
Let $I_{f}$ denote the saturation of the ideal $J_{f}$ with respect to the maximal ideal ${\bf m}=(x,y,z)$ in $S$ and consider the graded Jacobian module of $f$ defined by
$N(f)=I_{f}/J_{f}.$
We set $n(f)_{k}=\dim N(f)_{k}$ for any integer $k$ and introduce the freeness defect of the curve $C$ by the formula
$\nu(C)=\max_{j}\{n(f)_{j}\}$
as in [2]. Note that $C$ is free if and only if $N(f)=0$, see for instance [24], and hence in this case $\nu(C)=0$, and $C$ is nearly free if and only if $\nu(C)=1$, see [13]. To prove Theorem 1.2, we show in fact that $\tau(C)=\tau(d,r)_{max}-2$ implies $\nu(C)=2$ and then use the classification of such curves given in [15, Theorem 3.11]. This result says that a curve $C$ with $\nu(C)=2$ is either an MPOG-curve, or a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$, which is in fact a maximal Tjurina curve of type $(d,r)=(2r-1,r)$ as defined in [16], see also Section 2 below.
The Briançon-type polynomials are non-homogeneous polynomials $h$ in the ring of polynomials $R=\mathbb{C}[x,y]$, such that the associated mapping $h:\mathbb{C}^{2}\to\mathbb{C}$ has all the fibers $h^{-1}(t)$ smooth and irreducible, and yet $h$ is not obtained from a linear form in $R$ by composition with a polynomial automorphism of $\mathbb{C}^{2}$. A series of such polynomials have been constructed in [3], the simplest of them being two polynomials of degree 10 that we recall now.
Let $s=xy+1$, $p=xs+1$ and $u=s^{2}+y$. Using this notation, we define
(1.3) $g=p^{2}u-\frac{5}{3}ps-\frac{1}{3}s\text{ and }g^{\prime}=p^{2}u-\frac{7}{9}ps+\frac{1}{9}s.$
The polynomial $g$ up-to a change of coordinates $(x,y)\mapsto(-x,-y)$ was considered in an unpublished manuscript by J. Briançon, see [3, 4]. We recall that for a polynomial mapping $h:\mathbb{C}^{2}\to\mathbb{C}$ the set of atypical values is the minimal set $A(h)\subset\mathbb{C}$ such that $h$ induces a locally trivial topological fibration over the set $\mathbb{C}\setminus A(h)\cup C(h)$, where $C(h)$ denotes the set of critical values of $h$. The atypical values for a function $h:\mathbb{C}^{2}\to\mathbb{C}$ come from jumping of the Milnor numbers at infinity, see [5, 18, 20].
The polynomials $g$ and $g^{\prime}$ have the following very interesting properties, see [3].
Theorem 1.3.
  1. (1)
    The polynomials $g$ and $g^{\prime}$ have no critical points, that is all the fibers of $g$ and $g^{\prime}$ are smooth affine curves. In other words, the sets $C(g)$ and $C(g^{\prime})$ are empty.
  2. (2)
    All the fibers of the polynomials $g$ and $g^{\prime}$ are irreducible affine plane curves.
  3. (3)
    The set of atypical values for $g$ is $A(g)=\{0,b\}$, where $b=-\frac{16}{9}.$
  4. (4)
    The set of atypical values for $g^{\prime}$ is $A(g^{\prime})=\{0,b^{\prime}\}$, where $b^{\prime}=-\frac{64}{81}.$
  5. (5)
    The generic fiber $F_{1}=g^{-1}(1)$ of the polynomial $g$ and $F_{1}^{\prime}=g^{\prime-1}(1)$ of the polynomial $g$ have their first Betti numbers given by $b_{1}(F_{1})=b_{1}(F_{1}^{\prime})=4$. In particular, the polynomials $g$ and $g^{\prime}$ are not obtained from linear forms by composition with a polynomial automorphism of $\mathbb{C}^{2}$.
When $h:\mathbb{C}^{2}\to\mathbb{C}$ is obtained from a linear form in $R$ by composition with a polynomial automorphism of $\mathbb{C}^{2}$, the freeness properties of the projective closure $D_{t}$ of any fiber $h^{-1}(t)$ are discussed in [11, Theorem 1.5]. It is natural to ask what happens when we replace such a polynomial $h$ by a Briançon-type polynomial.
For the polynomial $g$, we consider three associated plane curves $C_{t}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $F_{t}=g^{-1}(t)$ for $t=0$, $t=b$ and $t=1$. The first two fibers correspond to the atypical values in $A(g)$, and the third one is a general fiber of $g$. For the convenience of the reader, the defining equations for these curves, obtained by homogeneization starting with the first equation in (1.3), are given in (3.1) and (3.2). Similarly, for the polynomial $g^{\prime}$, we consider three associated plane curves $C_{t}^{\prime}$ in $\mathbb{P}^{2}$, which are the projective closure of the affine fibers $g^{\prime-1}(t)$ for $t=0$, $t=b^{\prime}$ and $t=1$. The corresponding equations are given in (3.3) and (3.4). Here are our results about the freeness properties of these projective plane curves.
Theorem 1.4.
The geometric genera of the curves $C_{t}$ are given by
$g(C_{1})=1\text{ and }g(C_{0})=g(C_{b})=0.$
Moreover, these curves of degree $10$ enjoy the following freeness properties.
  1. (1)
    The curves $C_{0}$ and $C_{1}$ are $MPOG$-curves with exponents $(5,5,6)$ and $\tau(C_{0})=\tau(C_{1})=59$.
  2. (2)
    The curve $C_{b}$ is a free curve with exponents $(4,5)$ and $\tau(C_{b})=61$.
Theorem 1.5.
The geometric genera of the curves $C_{t}^{\prime}$ are given by
$g(C_{1}^{\prime})=1\text{ and }g(C_{0}^{\prime})=g(C_{b^{\prime}}^{\prime})=0.$
Moreover, these curves of degree $10$ enjoy the following freeness properties.
  1. (1)
    The curve $C_{1}^{\prime}$ is an $MPOG$-curve with exponents $(5,5,6)$ and $\tau(C_{1}^{\prime})=59$.
  2. (2)
    The curves $C_{0}^{\prime}$ and $C_{b^{\prime}}^{\prime}$ are nearly free curves with exponents $(5,5)$ and $\tau(C_{0}^{\prime})=\tau(C_{b^{\prime}}^{\prime})=60$.
Remark 1.6.
The claims in Theorem 1.4 and Theorem 1.5 about the genera of the curves $C_{t}$ and $C_{t}^{\prime}$ are already proved in [3]. We state them here for the following reason. Among the six curves considered in Theorem 1.4 and Theorem 1.5, the curve $C_{b}$ is rational cuspidal and the curves $C_{0}^{\prime}$ and $C_{b^{\prime}}^{\prime}$ are rational nearly cuspidal curves, that is all their singularities are irreducible except one which has 2 branches, see Remarks 3.1 and 3.2. Hence the claims above are compatible with fact that a rational cuspidal curve of even degree is either free or nearly free, see [13], and a rational nearly cuspidal curve $C$ of even degree is free, nearly free or an $MPOG$-curve, see [15, 13]. The curve $C_{0}$, which is not cuspidal, shows moreover that the hypothesis of being cuspidal is necessary in general for the first result to hold, even if there are exceptions, as shown by the curves $C_{0}^{\prime}$ and $C_{b^{\prime}}^{\prime}$ which are not cuspidal. The curves $C_{1}$ and $C_{1}^{\prime}$ are both $MPOG$-curves, even though they are not rational.
Since the line at infinity $L_{z}:z=0$ intersects the curves $C_{t}$ only in two points, the general principle explained in [12] tells us that the curves $C_{t}\cup L_{z}$ (resp. $C_{t}^{\prime}\cup L_{z}$) of degree $11$ may also enjoy (sometimes better) freeness properties. See also Step 3 in the proof of Theorem 1.5 in [11] for such a situation. We consider also the curve arrangements $C_{0}\cup C_{b}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}$) and $C_{0}\cup C_{b}\cup L_{z}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$. Here are the corresponding results.
Theorem 1.7.
  1. (1)
    The curves $C_{0}\cup L_{z}$ and $C_{1}\cup L_{z}$ are nearly free with exponents $(5,6)$ and $\tau(C_{0}\cup L_{z})=\tau(C_{1}\cup L_{z})=74$.
  2. (2)
    The curve $C_{b}\cup L_{z}$ is free with exponents $(4,6)$ and $\tau(C_{b}\cup L_{z})=76$.
  3. (3)
    The curve $C_{0}\cup C_{b}$ (resp. $C_{0}\cup C_{b}\cup L_{z}$) is free with exponents $(9,10)$ and $\tau(C_{0}\cup C_{b})=271$, (resp. exponents $(9,11)$) and $\tau(C_{0}\cup C_{b}\cup L_{z})=301$.
Theorem 1.8.
  1. (1)
    The curves $C_{0}^{\prime}\cup L_{z}$ and $C_{1}^{\prime}\cup L_{z}$ are nearly free with exponents $(5,6)$ and $\tau(C_{0}^{\prime}\cup L_{z})=\tau(C_{1}^{\prime}\cup L_{z})=74$.
  2. (2)
    The curve $C^{\prime}_{b^{\prime}}\cup L_{z}$ is free with exponents $(5,5)$ and $\tau(C^{\prime}_{b^{\prime}}\cup L_{z})=75$.
  3. (3)
    The curve $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$) is nearly free with exponents $(9,11)$ and $\tau(C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}})=270$, (resp. exponents $(9,12)$ and $\tau(C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z})=300$).
In the pencils of curves ${\alpha}C_{0}+{\beta}C_{b}$ (resp. ${\alpha}C_{0}^{\prime}+{\beta}C_{b^{\prime}}^{\prime}$) there are 3 exceptional members (i.e., members more singular than the general member), namely $C_{0},C_{b}$ and $10L_{z}$ (resp. $C_{0}^{\prime}$, $C_{b^{\prime}}^{\prime}$. and $10L_{z}$). Therefore the freeness of the curve $C_{0}\cup C_{b}\cup L_{z}$ (resp. $C_{0}^{\prime}\cup C^{\prime}_{b^{\prime}}\cup L_{z}$) can be related to the results in [7], see also [9, 26].
Note that the lines $L_{x}:x=0$ and $L_{y}:y=0$ satisfy
$|C_{0}\cap L_{x}|=|C_{0}^{\prime}\cap L_{x}|=2\text{ and }|C_{0}\cap L_{y}|=|C_{0}^{\prime}\cap L_{y}|=3,$
hence they are also high order inflectional lines as for the curves $C_{0}$ and $C_{0}^{\prime}$. Recalling our constructions in [12], it is natural to ask what happens if we add these lines to $C_{0}$ and to $C_{0}^{\prime}$. By adding the lines $L_{x}$ and $L_{y}$ to the curve $C_{0}$ we get new examples of maximal Tjurina curves of type $(d,r)=(2r-1,r)$ as defined above. The complete result is the following.
Theorem 1.9.
The curve $C_{0}\cup L_{x}$ is nearly free with exponents $(5,6)$, the curve $C_{0}\cup L_{y}$ is maximal Tjurina of type $(11,6)$ and the curve $C_{0}\cup L_{x}\cup L_{y}\cup L_{z}$ is maximal Tjurina of type $(13,7)$.
The curve $C_{0}^{\prime}\cup L_{x}$ is free with exponents $(5,5)$, the curve $C_{0}^{\prime}\cup L_{y}$ is nearly free with exponents $(5,6)$ and the curve $C_{0}^{\prime}\cup L_{x}\cup L_{y}\cup L_{z}$ is nearly free with exponents $(6,7)$.
Using the results in [3], one can construct other new curves of higher degree, which are either free, or nearly free or MPOG-curves, see Section 4 below. This leads us to ask the natural question if such a result holds for any Briançon-type polynomials.
Question 1.10.
Do the projective closures of the general and the special (atypical) fibers of any Briançon-type polynomial enjoy some kind of freeness property, i.e., do the corresponding exponents have some special properties ?
It seems that this question is related to the very challenging question of finding numerical restrictions for Briançon-type polynomials, involving the number of special fibers and their geometry.
2. A new characterization of $MPOG$-curves
Recall that the total Tjurina number $\tau(C)$ of a reduced plane curve $C:f=0$ is just the degree of its Jacobian ideal $J_{f}$, or equivalently the sum of the Tjurina numbers of all the singularities of $C$. The following result, due to du Plessis and Wall, see [17, Theorem 3.2] as well as [19] for an alternative approach, gives bounds on this Tjurina number $\tau(C)$.
Theorem 2.1.
For positive integers $d$ and $r$, define two new integers by
$\tau(d,r)_{min}=(d-1)(d-r-1)\text{ and }\tau(d,r)_{max}=(d-1)^{2}-r(d-r-1).$
Then, if $C:f=0$ is a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$, one has
$\tau(d,r)_{min}\leq\tau(C)\leq\tau(d,r)_{max}.$
Moreover, for $r={\rm mdr}(f)\geq d/2$, the stronger inequality $\tau(C)\leq\tau^{\prime}(d,r)_{max}$ holds, where
$\tau(d,r)^{\prime}_{max}=\tau(d,r)_{max}-\binom{2r+2-d}{2}.$
A degree $d$ curve $C:f=0$ satisfying $r={\rm mdr}(f)\geq d/2$ and $\tau(C)=\tau^{\prime}(d,r)_{max}$ is called a maximal Tjurina curve of type $(d,r)$, see [16].
At the end of the proof of Theorem 2.1, in [17], the authors state the following very interesting consequence (of the proof, not of the statement) of Theorem 2.1.
Corollary 2.2.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$. One has
$\tau(C)=\tau(d,r)_{max}$
if and only if $C:f=0$ is a free curve, and then $r<d/2$.
In the paper [8], the first author has given an alternative proof of Corollary 2.2 and has shown that a similar property holds for nearly free curves.
Corollary 2.3.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$. One has
$\tau(C)=\tau(d,r)_{max}-1$
if and only if $C:f=0$ is a nearly free curve, and then $r\leq d/2$.
2.4. The proof of Theorem 1.2
If we assume that $C$ is an $MPOG$-curve, then the equality for $\tau(C)$ follows from [15, Proposition 2.1, (4)].
Conversely, assume now that $\tau(C)=\tau(d,r)_{max}-2.$ If $r>d/2$, then Theorem 2.1 implies that
$\tau(C)\leq\tau(d,r)_{max}^{\prime}\leq\tau(d,r)_{max}-\binom{3}{2}=\tau(d,r)_{max}-3,$
a contradiction. It follows that $r\leq d/2$. Now [10, Theorem 1.2] implies that
$\nu(C)=\tau(d,r)_{max}-\tau(C)=2,$
when $r<d/2$. In the case $r=d/2$, we have the following equalities, as in the discussion after [10, Theorem 1.2].
$\nu(C)=3r^{2}-3r+1-\tau(C)=2,$
since $\tau(C)=\tau(d,r)_{max}-2=3r^{3}-3r-1.$ Therefore the assumption $\tau(C)=\tau(d,r)_{max}-2$ implies in all cases that $\nu(C)=2$. Now we apply [15, Theorem 3.11] and conclude that $C$ is an $MPOG$-curve.
Remark 2.5.
Let $C:f=0$ be a reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $r={\rm mdr}(f)$. Then the equality
$\tau(C)=\tau(d,r)_{max}-3$
does not determine uniquely the exponents of $C$. Indeed, such a curve may be a plus-one generated curve with exponents $(r,d-r,d-r+2)$, a 4-syzygy curve with exponents $(r,r,r,r)$ and degree $d=2r-1$ as in [15, Theorem 3.11], or maybe even other types of curves.
Remark 2.6.
The free and nearly free curves have strong relations with the rational cuspidal plane curves, see [13, 14]. The $MPOG$-curves are related with the rational nearly cuspidal curves, i.e. rational curves whose singularities are all irreducible except one singularity, which has two branches, see [15, Theorem 1.3, Conjecture 1.4, Corollary 5.8]. See also [21] for an alternative view on nearly free curves.
3. On the freeness of the curves associated to Briançon-type polynomials of degree 10
Here are the defining equations for the curves $C_{t}$ in $\mathbb{P}^{2}$, obtained by homogeneization starting with the first equation in (1.3).
(3.1) $C_{0}:f_{0}=x^{6}y^{4}+4x^{5}y^{3}z^{2}+3x^{4}y^{3}z^{3}+6x^{4}y^{2}z^{4}+\frac{19}{3}x^{3}y^{2}z^{5}+4x^{3}yz^{6}+3x^{2}y^{2}z^{6}+$
$+\frac{11}{3}x^{2}yz^{7}+x^{2}z^{8}+2xyz^{8}+\frac{1}{3}xz^{9}+yz^{9}-z^{10}=0,$
(3.2) $C_{b}:f_{b}=f_{0}+\frac{16}{9}z^{10}=0\text{ and }C_{1}:f_{1}=f_{0}-z^{10}=0.$
Remark 3.1.
It is clear that all the curves $C_{t}$ have two points situated on the line at infinity $L:z=0$, namely the points $p=(1:0:0)$ and $q=(0:1:0)$. It is easy to check using SINGULAR [6] that one has
$\mu(C_{0},q)=\mu(C_{b},q)=\mu(C_{1},q)=42$
and
$\tau(C_{0},q)=\tau(C_{b},q)=\tau(C_{1},q)=35.$
Hence at the point $q$ the family of plane curve singularities $(C_{t},q)$ is $\mu$-constant and $\tau$-constant. Moreover, all these curve singularities have just one branch, i.e., they are irreducible. On the other hand we have
$\mu(C_{0},p)=28,\ \mu(C_{b},p)=30\text{ and }\mu(C_{1},p)=27.$
The Tjurina numbers also vary, as we have
$\tau(C_{0},p)=\tau(C_{1},p)=24\text{ and }\tau(C_{b},p)=26.$
It is interesting that the total Tjurina number of the exceptional curve $C_{0}$ coincides with the total Tjurina number of the generic curve $C_{1}$. Indeed, recall that the exceptional fibers $F_{t}$ of a polynomial mapping $g:\mathbb{C}^{2}\to\mathbb{C}$ are detected by the jump in the total Milnor number of the compactifications $C_{t}$, see [20]. The number of branches $r(C_{t},p)$ for the singularity $(C_{t},p)$ is given by the following
$r(C_{0},p)=3,\ r(C_{b},p)=1\text{ and }r(C_{1},1)=2.$
In fact, the claims about the Milnor numbers $\mu(C_{t},p)$ and $\mu(C_{t},q)$, and about the number of branches $r(C_{t},p)$ and $r(C_{t},q)$ are also proven in a different way in [3].
Here are the defining equations for the curves $C^{\prime}_{t}$, obtained by homogeneization starting with the second equation in (1.3).
(3.3) $C_{0}^{\prime}:f_{0}^{\prime}=x^{6}y^{4}+4x^{5}y^{3}z^{2}+3x^{4}y^{3}z^{3}+6x^{4}y^{2}z^{4}+\frac{65}{9}x^{3}y^{2}z^{5}+4x^{3}yz^{6}+3x^{2}y^{2}z^{6}+$
$+\frac{49}{9}x^{2}yz^{7}+x^{2}z^{8}+\frac{10}{3}xyz^{8}+\frac{11}{9}xz^{9}+yz^{9}+\frac{1}{3}z^{10}=0,$
(3.4) $C_{b^{\prime}}^{\prime}:f_{b^{\prime}}^{\prime}=f_{0}^{\prime}+\frac{64}{81}z^{10}=0\text{ and }C_{1}^{\prime}:f_{1}^{\prime}=f_{0}^{\prime}-z^{10}=0.$
Remark 3.2.
It is clear that all the curves $C_{t}^{\prime}$ have two points situated on the line at infinity $L:z=0$, namely the points $p=(1:0:0)$ and $q=(0:1:0)$. It is easy to check using SINGULAR that one has
$\mu(C_{0}^{\prime},q)=\mu(C_{b^{\prime}}^{\prime},q)=\mu(C_{1}^{\prime},q)=42$
and
$\tau(C_{0}^{\prime},q)=\tau(C_{b^{\prime}}^{\prime},q)=\tau(C_{1}^{\prime},q)=35.$
Hence at the point $q$ the family of plane curve singularities $(C_{t}^{\prime},q)$ is $\mu$-constant and $\tau$-constant. Moreover, all these curve singularities have just one branch. On the other hand we have
$\mu(C_{0}^{\prime},p)=\mu(C_{b^{\prime}}^{\prime},p)=29\text{ and }\mu(C_{1}^{\prime},p)=27.$
The Tjurina numbers also vary, as we have
$\tau(C_{0}^{\prime},p)=\tau(C_{b^{\prime}}^{\prime},p)=25\text{ and }\tau(C_{1}^{\prime},p)=24.$
Moreover, all these curve singularities at $p$ have two branches. These results, except the computation of the Tjurina numbers, are also obtained in an alternative way in [3].
3.3. The proofs of Theorems 1.4, 1.5, 1.7, 1.8 and 1.9
All these proofs involve the use of some computations done using a computer algebra software, for instance SINGULAR [6]. One can proceed in several ways, as follows.
  1. (1)
    Use the defining equations given in (3.1), (3.2), (3.3) and (3.4) and list a minimal set of generators for the corresponding module of Jacobian syzygies. Then the values for the total Tjurina numbers can be obtained using Corollaries 2.2 and 2.3, or our new Theorem 1.2.
  2. (2)
    Use the defining equations given in (3.1), (3.2), (3.3) and (3.4) to compute the minimal degree of a Jacobian syzygy in each case, compute the total Tjurina numbers as in Remarks 3.1 and 3.2 and then use Corollaries 2.2 and 2.3, or Theorem 1.2 to decide whether the curve is free, or nearly free, or an MPOG-curve.
One may use both approaches to check the results given by the computer algebra software.
4. On higher degree Briançon-type polynomials
The authors in [3] have constructed (several) Briançon-type polynomials $g_{n}$ for any degree $d_{n}=6n+4$, the polynomials $g$ and $g^{\prime}$ in (1.3) being the polynomials corresponding to $n=1$. They are given by the formula
(4.1) $g_{n}=p^{2n}u+s\left(\sum_{j=0}^{n}a_{j}p^{j}+\sum_{j=n+1}^{2n-1}p^{j}\right),$
where $s,p,u$ are as in (1.3) and the coefficients $a_{0},\ldots,a_{n}$ take specified values in $\mathbb{C}$. For $n=2$, there are 3 such polynomials constructed in [3], the simplest good choice (corresponding to Case 1 in [3]) of the triple of coefficients being
(4.2) $a_{0}=-\frac{1}{5},\ a_{1}=-\frac{3}{5}\text{ and }a_{2}=-\frac{11}{5}.$
The corresponding polynomial, to be denoted by $g^{\prime\prime}$, has 2 atypical values, 00 and
$b^{\prime\prime}=-(a_{2}-1)^{2}/4=-\frac{64}{25}.$
The corresponding curves in $\mathbb{P}^{2}$, denoted by $C_{0}^{\prime\prime}$, $C_{b^{\prime\prime}}^{\prime\prime}$ and $C_{1}^{\prime\prime}$ have similar properties to the curves $C_{t}$ and $C_{t}^{\prime}$ discussed above. The equations for these curves are the following.
$C_{0}^{\prime\prime}:f_{0}^{\prime\prime}=x^{10}y^{6}+6x^{9}y^{5}z^{2}+5x^{8}y^{5}z^{3}+15x^{8}y^{4}z^{4}+25x^{7}y^{4}z^{5}+20x^{7}y^{3}z^{6}+10x^{6}y^{4}z^{6}+$
$+50x^{6}y^{3}z^{7}+15x^{6}y^{2}z^{8}+184/5x^{5}y^{3}z^{8}+50x^{5}y^{2}z^{9}+10x^{4}y^{3}z^{9}+6x^{5}yz^{10}+$
$+252/5x^{4}y^{2}z^{10}+25x^{4}yz^{11}+22x^{3}y^{2}z^{11}+x^{4}z^{12}+152/5x^{3}yz^{12}+5x^{2}y^{2}z^{12}+$
$+5x^{3}z^{13}+14x^{2}yz^{13}+34/5x^{2}z^{14}+4xyz^{14}+2xz^{15}+yz^{15}-z^{16}=0,$
$C_{1}^{\prime\prime}:f_{1}^{\prime\prime}=f_{0}^{\prime\prime}-z^{16}=0\text{ and }C_{b^{\prime\prime}}:f_{b^{\prime\prime}}^{\prime\prime}=f_{0}+\frac{64}{25}z^{16}=0.$
It was shown in [3] that $g(C_{0}^{\prime\prime})=g(C_{b^{\prime\prime}}^{\prime\prime})=0$, $g(C_{1}^{\prime\prime})=2$ and all the curves $C_{t}^{\prime\prime}$ have two points on the line at infinity $L$. The interested reader can check, using one of the approaches $(1)$ or $(2)$ explained above at the end of Section 3, the following result.
Theorem 4.1.
  1. (1)
    The curves $C_{0}^{\prime\prime}$ and $C_{1}^{\prime\prime}$ are MPOG-curves with exponents $(8,8,9)$ and $\tau(C_{0}^{\prime\prime})=\tau(C_{1}^{\prime\prime})=167$. The curve $C_{b^{\prime\prime}}^{\prime\prime}$ is free with exponents $(7,8)$ and $\tau(C_{b^{\prime\prime}}^{\prime\prime})=169$.
  2. (2)
    The curve $C_{b^{\prime\prime}}^{\prime\prime}\cup L$ is free with exponents $(7,9)$ and $\tau(C_{b^{\prime\prime}}^{\prime\prime}\cup L)=193$ and the curves $C_{0}^{\prime\prime}\cup L$ and $C_{1}^{\prime\prime}\cup L$ are nearly free with exponents $(8,9)$ and $\tau(C_{0}^{\prime\prime}\cup L)=\tau(C_{1}^{\prime\prime}\cup L)=191$.
  3. (3)
    The curve $C_{0}^{\prime\prime}\cup C_{b^{\prime\prime}}^{\prime\prime}$ is free with exponents $(15,16)$ and $\tau(C_{0}^{\prime\prime}\cup C_{b^{\prime\prime}}^{\prime\prime})=721$ and the curve $C_{0}^{\prime\prime}\cup C_{b^{\prime\prime}}^{\prime\prime}\cup L$ is free with exponents $(15,17)$ and $\tau(C_{0}^{\prime\prime}\cup C_{b^{\prime\prime}}^{\prime\prime})=769$.
For each $n\geq 2$, the authors in [3] have shown the existence of a specific choice of the coefficients in (4.1), corresponding to the Case $n+1$ in [3]. Explicit formulas for the polynomial $g_{n}^{\prime}$ obtained from this choice are given in [3] for the cases $2\leq n\leq 4$, which involve rational numbers with large denominators, see Remark 4.3.
For this polynomial $g_{n}^{\prime}$, the projective closure $C_{n,0}^{\prime}$ of the affine fiber $g_{n}^{{}^{\prime}-1}(0)$ is a rational curve, having 3 branches at infinity by [3, Figure 7]. Since $C_{n,0}^{\prime}$ has 2 points at infinity, it follows that $C_{n,0}^{\prime}$ is a nearly cuspidal rational curve, as defined in Remark 2.6 above. Using the proof of [15, Theorem 1.3] and the statement of [15, Theorem 3.11], we obtained the following.
Theorem 4.2.
For any $n\geq 2$, the curve $C_{n,0}^{\prime}$ coming from a special fiber of the Briançon-type polynomial $g_{n}^{\prime}$ is either free, or nearly free or an MPOG-curve.
Remark 4.3.
A direct computation with SINGULAR shows that $C_{2,0}^{\prime}$ is a free curve of degree 16 with exponents $(7,8)$, $C_{3,0}^{\prime}$ is a free curve of degree 22 with exponents $(7,14)$ and $C_{4,0}^{\prime}$ is a free curve of degree 28 with exponents $(7,20)$. The choice of the coefficients $a_{j}$’s, that gives rise to the polynomial $g_{n}^{\prime}$ are listed in [3] as follows. For $n=2$ the coefficients are
$a_{0}=-\frac{1}{125},\ a_{1}=\frac{17}{125}\text{ and }a_{2}=-\frac{91}{125},$
for $n=3$ the coefficients are
$a_{0}=\frac{1}{2401},\ a_{1}=-\frac{31}{2401},\ a_{2}=\frac{353}{2401}\text{ and }a_{3}=-\frac{1695}{2401}$
and the coefficients for the case $n=4$ are the following
$a_{0}=-\frac{1}{N},\ a_{1}=\frac{49}{N},\ a_{2}=-\frac{951}{N},\ a_{3}=\frac{9049}{N}\text{ and }a_{4}=-\frac{40951}{N},$
where $N=59049$. We do not know whether the curves $C_{n,0}^{\prime}$ are free for any $n\geq 5$.